OpenStudy (anonymous):

Fool's problem of the day, Today's problems are to find the next term(s) of some given series (a) \( 7,9,40,74,1526, ? \) \( 5436 \quad\quad 6486 \quad\quad5486 \quad\quad3456 \) (b) \( 1,3,8,18,? \) \( 26 \quad\quad 30 \quad\quad 32 \quad\quad 38 \) (c) \( Z, X, S,I,R,R,?,? \) \( JK \quad\quad JI \quad\quad KM \quad\quad GI \) Please post the solution instead of just the answer. PS: Only the first one is relatively easy/well-known.

OpenStudy (lgbasallote):

fool never fails to disappoint..always nearly unsolveable =)))

OpenStudy (anonymous):

for (b), 35?

OpenStudy (anonymous):

@CoCoTsoi: Sorry, 35 is the not the right answer.

OpenStudy (anonymous):

Happy to learn from failure :D

OpenStudy (lgbasallote):

these are not like college exams that have common patterns...the progressions seems inconsistent..but still seems not...

OpenStudy (anonymous):

it is quite hard to know whether we think it correctly coz the sequence given can deduct numerous sequence that are different

OpenStudy (anonymous):

for (b), 68?

OpenStudy (anonymous):

@CoCoTsoi: I have added option now :)

OpenStudy (anonymous):

that's excellent!

OpenStudy (anonymous):

(b) 26

OpenStudy (anonymous):

Please post solution instead of just an answer.

OpenStudy (lgbasallote):

go Coco! be the first person to finally answer Fool's question for the day! represent OpenStudy!!! hahaha =)))

OpenStudy (experimentx):

for b) tn = tn+(n^2+1) works out ... but there not in the answer

OpenStudy (lgbasallote):

i dont know what these mean..goodluck with these =)))) <insert yaoming meme>

OpenStudy (anonymous):

Help, I am confused by myself. Wait a little bit longer plz

OpenStudy (experimentx):

tn = t(n-1)+((n-1)^2+1)

OpenStudy (anonymous):

No one has posted the correct answer to any one of them.

OpenStudy (experimentx):

hark working very very working ...!!

OpenStudy (lgbasallote):

and you're wondering....why? lol =))) hey fool..has anyone answered one of your Fool's problem of the day?

OpenStudy (anonymous):

@lgbasallote: Yes many did.

OpenStudy (experimentx):

man .. i was wondering who what order would give such plain and sharp rise

OpenStudy (lgbasallote):

wow..they must feel proud...lemme guess...ishaan, diyadiya, saifoo, bahrom, those folks right?

OpenStudy (callisto):

For a, is it 5436? I dunno what i'm doing.. third term = (1st term)^2 - second term = 7^2 -9 4th term = 2nd term ^2 - first term 5th term = 3rd term^2 - 4th term So i'm guessing 6th term = 4th term^2 - 3rd term @Callisto, it's not correct :(

OpenStudy (lgbasallote):

why did callisto tag herself? i wonder....hahaha

OpenStudy (anonymous):

for b) 32?

OpenStudy (anonymous):

I would have chosen 35 for b as well...

OpenStudy (karatechopper):

i think, a) 5436

OpenStudy (mani_jha):

For the first one, I thought the same way Callisto did. For the second one, I guess 26(The terms are 1^2-0,2^2-1,3^2-1,4^2+2, 5^2+1) @order and @anonymoustwo44, Please post your explanations.

OpenStudy (callisto):

@Mani_Jha but then that's doesn't look like a series for b

OpenStudy (callisto):

*delete ' 's ' please*

OpenStudy (anonymous):

I thought (b) 1,3,8,18,35 since, there's a difference of 2, 5, 10, 17, since 1-2 = 1, 2-5=3, 5-10=5, so, 10-x = 7 = 17,

OpenStudy (mani_jha):

I know. That's why I said I 'guess' :P

OpenStudy (anonymous):

callisto is right for the first one! Congratz :D @anonymoustwo44: Solution please.

OpenStudy (anonymous):

Why isn't 42 in the options? hehe

OpenStudy (callisto):

i wish there could be a 31 in the option :(

OpenStudy (anonymous):

maybe this,\[a_0 = 1\] \[a_{n} = n^2 + 1 + a_{n-1}, n \in \mathbb{N}\]

OpenStudy (anonymous):

but 35 isn't the option

OpenStudy (anonymous):

I hate these questions

OpenStudy (anonymous):

for b) 2^0=1, 1+2^1=3, 3+2^2+2^0=6,8+2^1+2^3=18. 18+2^2+2^4 = 38?

OpenStudy (anonymous):

correction : 2^0=1, 1+2^1=3, 3+2^2+2^0=8,8+2^1+2^3=18, 18+2^2+2^4 = 38?

OpenStudy (anonymous):

Can you generalize your series?

OpenStudy (turingtest):

I seeing the answer to b) as 30 but am having difficulty explaining the first two terms

OpenStudy (turingtest):

1*3=3 3*3-1*1=8 8*3-2*3=18 18*3-3*8=30

OpenStudy (anonymous):

I guess some weird formula like a(n)=a(n-1)+2^(n-3)+2^(n-1), but I don't know if I can just not take into account the negative exponants :-S

OpenStudy (turingtest):

b) 1=1 1*3=3 3*3-1*1=8 8*3-2*3=18 18*3-3*8=30\[x_0=1\]\[x_n=3x_{n-1}-(n-1)x_{n-2}\]maybe, looks too simple though

OpenStudy (anonymous):

For 3rd it's GI

OpenStudy (turingtest):

how do you figure?

OpenStudy (turingtest):

I think I'm right for b) hehe

OpenStudy (anonymous):

It's relatively easy, just list out the alphabets, series follow \(a_n = n^2 + 1, n \in \mathbb{N}\) from the bottom.

OpenStudy (anonymous):

or, Z to A.

OpenStudy (turingtest):

I see :)

OpenStudy (anonymous):

a(n+1) = a(n-1)^2 - a(n - 2*(1 - n mod 2)), a(0)=7, a(1)=9

OpenStudy (anonymous):

Congratz! ishaan, you got the right answer of (3)

OpenStudy (anonymous):

For 2: Answer is 32 and hint is "Prime numbers"

OpenStudy (turingtest):

aww... I thought I had it for once! I do realize I had a problem with the fact that \(x_{-1}\) is not defined ok I'll keep staring at it...

OpenStudy (anonymous):

3+5 = 8, 7+11 = 18, 13+19 = 32, so it looks a lot like it's ((n-1)*2-2)th + ((n-1)*2-1)th primes, but that doesn't really explain 1 and 3, does it?

OpenStudy (inkyvoyd):

Are these polynomials?

OpenStudy (anonymous):

@inkyvoyd: Interpolation won't help.

OpenStudy (inkyvoyd):

*I'm out*

OpenStudy (anonymous):

I give up :/ ...even after hint I wasn't able to do it :( ...Post the logic FoolForMath if others think so too, I know I am gonna feel stupid after you post the solution but feeling stupid for an hour is better than feeling hopeless for several hours.

OpenStudy (turingtest):

1+2=3 3+5=8 7+11=18 13+17=30 with that way of summing pairs of consecutive primes I got 30 again lol it looks like you want me to sum 19 and 13, but why...?

OpenStudy (anonymous):

I agree with Ishaan, I'm really curious about this

OpenStudy (anonymous):

Solution of (b): Here, in this series we are omitting prime numbers in an increasing order. Between 1 and 3 we are omitting '2' i.e., one prime number. Between 3 and 8 we are omitting 5 and 7 i.e. two prime numbers. Similarly between 18 and missing number we have to omit 4 prime numbers. The next four prime numbers after 18 are 19,23,29,31. Hence,32 is the missing term.

OpenStudy (anonymous):

That's it? A very unusual series.

OpenStudy (anonymous):

I agree:)